### v1

# Common and Rare Fundus Diseases Identification Using Vision-Language Foundation Model with Knowledge of Over 400 Diseases

Abstract
The current retinal artificial intelligence models were trained using data with a limited category of diseases and limited knowledge. In this paper, we present a retinal vision-language foundation model (RetiZero) with knowledge of over 400 fundus diseases. Specifically, we collected 341,896 fundus images paired with text descriptions from 29 publicly available datasets, 180 ophthalmic books, and online resources, encompassing over 400 fundus diseases across multiple countries and ethnicities. RetiZero achieved outstanding performance across various downstream tasks, including zero-shot retinal disease recognition, image-to-image retrieval, internal domain and cross-domain retinal disease classification, and few-shot fine-tuning. Specially, in the zero-shot scenario, RetiZero achieved a Top5 score of 0.8430 and 0.7561 on 15 and 52 fundus diseases respectively. In the image-retrieval task, RetiZero achieved a Top5 score of 0.9500 and 0.8860 on 15 and 52 retinal diseases respectively. Furthermore, clinical evaluations by ophthalmology experts from different countries demonstrate that RetiZero can achieve performance comparable to experienced ophthalmologists using zero-shot and image retrieval methods without requiring model retraining. These capabilities of retinal disease identification strengthen our RetiZero foundation model in clinical implementation.
1 Introduction
Blindness and visual impairment represent a significant disease burden globally, impacting millions of individuals. Detection and timely treatment of ocular conditions, such as retinal and optic nerve diseases, are crucial for reducing severe and permanent damage. However, the insufficient availability of ophthalmic medical resources severely limits the prompt screening and treatment of retinal diseases. In recent years, artificial intelligence (AI)-based retinal disease screening systems have alleviated the workload on healthcare professionals to some extent, providing crucial technological support for the timely screening of retinal disease patients and referring for treatment. Nevertheless, most previous AI-based methods were customized for specific diseases, such as diabetic retinopathy (DR) [1, 2], glaucoma [3, 4], and retinopathy of prematurity (ROP) [5, 6]. Although several methods have been proposed for simultaneously screening multiple retinal diseases with promising performance [7, 8], most current AI models for ocular disease screening were trained on task-specific datasets, leading to inevitable errors in prediction when new data (ex., images acquired by different camera) or changes in tasks (ex., introducing new or rare categories). Furthermore, due to limited healthcare resources and the varying prevalence of retinal disease, collecting comprehensive datasets covering all retinal abnormalities is time-consuming and challenging. Consequently, most AI models were trained on limited data and disease categories, restricting their feature representation. Applying these models to different real-world settings or tasks requires extensive retraining with large datasets. Moreover, data quality and labeling issues further limit the widespread adoption of AI models in ophthalmic clinical settings.
Driven by abundant of big data and robust computing hardware, large foundation models (LFMs) have excelled in computer vision tasks [9, 10]. Pre-trained on massive datasets, LFMs provide rich feature support for downstream tasks like object detection [11], few-shot recognition [12], and zero-shot [13], etc. The first ophthalmic LFM, RETFound [14], introduced in 2023, was trained on large, unannotated retinal images using masked autoencoder (MAE) framework [15]. It provides rich feature support and improves the performance of downstream tasks, including internal domain and cross-domain retinal disease classification, few-shot learning, and prediction of systemic diseases. However, such an approach can hinder the model’s capacity to align feature information with labels in downstream tasks. In contrast, the Foundation LAnguage-Image model of the Retina (FLAIR), a Contrastive Language-Image Pre-training (CLIP) -based LFMs enhance feature representation by aligning text descriptions with image features, improving feature-label alignment but struggling with complex semantic features in medical imaging [10]. Current LFMs for ophthalmic imaging are pre-trained on extensive yet categorically limited datasets. Therefore, developing LFMs with comprehensive ophthalmic disease knowledge is crucial for representing complex retinal features and enhancing downstream task performance. However, collecting massive and diverse ophthalmic data that covers a wide range of retinal diseases for pretraining remains a significant challenge.
To address these problems and challenges, we collected 341,896 fundus images paired with text descriptions from 29 publicly available datasets (containing 303,124 fundus images with labels), 180 ophthalmic books (2,3328 fundus images with text descriptions), and online resources (15,544 fundus image with text descriptions), encompassing over 400 retinal and optic nerve diseases across multiple countries/regions and ethnicities. As shown in Figure  1, RetiZero is based on a contrastive vision-language pretraining framework that integrates MAE-based pretraining knowledge and low-rank training methods. Moreover, we introduced an uncertainty vision-language feature calibration method using Dirichlet reparameterization within the contrastive vision-language pretraining framework, to further better align vision-language features in the high-dimensional embedding space. Consequently, RetiZero achieved superior performance in various downstream tasks, including zero-shot fundus disease recognition, image-to-image fundus disease retrieval, internal domain retinal disease identification, few-shot fine-tuning, and cross-domain fundus disease identification.
2 Results
2.1 Zero-shot fundus disease recognition
The biggest advantage of RetiZero is the capability of zero-shot learning, which enables RetiZero to recognize fundus diseases using only textual prompts, without needing to retrain or fine-tune the model with labeled fundus images. As shown in Figure 2 (a), RetiZero achieved overall Top 1, Top 3, and Top 5 scores of 0.4421, 0.7024, 0.8404, respectively, for recognizing 15 common fundus diseases and normal condition of 30,089 fundus images (Eye-15 dataset), improved by 25.52%, 15.68%, and 15.61% over FLAIR (a recent VLM foundation model) [16], respectively. Furthermore, in the analysis of individual disease, RetiZero shows outstanding zero-shot capability in identifying most categories, especially for glaucoma (Top1: 0.7477, Top3: 0.9292, and Top5: 0.9718), Retinal Detachment (Top1: 0.7624, Top3: 0.9060, and Top5: 0.9628), and Retinitis Pigmentosa (Top1: 0.8190, Top3: 0.9533, and Top5: 0.9737) (Supplementary Figure 1). Additionally, to further validate RetiZero’s zero-shot capability in more challenging clinical scenarios, we collected a more demanding dataset named EYE-52. This dataset comprises 7,007 fundus images from various ophthalmology clinics, covering 52 fundus diseases, many of which are extremely rare in clinical practice. The incidence/prevalence of each category in the EYE-52 dataset was shown in Supplementary Table 3. As depicted in Figure 2 (b), RetiZero achieved overall Top1, Top3, and Top5 scores of 0.3595, 0.6259, and 0.7561, respectively, for recognizing the 52 types of fundus diseases in a zero-shot manner, providing superior performance compared FLAIR [16] (Top1: 0.0915, Top3: 0.2626, and Top5: 0.3398) and Random recognizing (Top1: 0.0294, Top3: 0.0882, and Top5: 0.1471). Furthermore,
Figure 1: Overview of the framework. Datasets for RetiZero pretraining: The RetiZero model was pre-trained using data from three primary sources: public datasets, ophthalmologic books, and online resources. We assembled a team of 10 ophthalmologists for manual data collection, with an additional two members dedicated to data cleaning tasks. This involved downloading images and corresponding labels from public datasets, extracting images and text descriptions from ophthalmic books, and downloading retinal diseases-relevant images and text descriptions from online resources. RetiZero combines the strengths of self-supervised learning based on the MAE architecture and contrastive learning from the CLIP architecture. Moreover, we introduce an uncertainty vision-language feature calibration method into the contrastive vision-language pretraining framework, to further calibrate visual-language features in the high-dimensional embedding space. Task I: Zero-shot fundus disease recognition. Task II: Fundus disease identification by image-to-image retrieval. Task III: Internal domain retinal disease identification. "Internal domain" means that we fine-tuned and tested the model using the data with similar feature distribution. Task IV: Few-shot fine-tuning. We fine-tuned the model using a few training samples to evaluate RetiZero’s performance in identifying fundus diseases with very limited training data. Task V: Cross-domain fundus disease identification. "Cross-domain" means that we fine-tuned and tested the model using the data with different feature distributions.
RetiZero demonstrated superior zero-shot performance, especially for recognizing some rare fundus diseases in clinical practice. For instance, RetiZero achieved Top 1, Top 3, and Top 5 scores of 0.6163, 0.7907, and 0.8605, respectively, for identifying Bietti Crystalline dystrophy. For the recognition of chorioretinal coloboma, the Top 1, Top 3, and Top 5 scores were 0.5085, 0.8079, and 0.9153, respectively (Supplementary Figure 2). Figure 2 (c) shows the Top5 prediction results provided by RetiZero and FLAIR for three rare disease samples. It can be seen that RetiZero’s Top5 predictions include the correct disease, further demonstrating the outstanding performance of RetiZero in screening rare diseases. More details on the rest of the 52 disease categories can be found in Supplementary Figure 2.
2.2 Fundus disease identification by image-to-image retrieval
As shown in task II of Figure 1, we sequentially randomly sampled one image from the dataset as the query image and used the remaining samples as the candidate pool, then we computed the similarity scores between the features extracted by the image encoder of RetiZero from the query image and all candidate images. Figure 2 (d) and Supplementary Figure 3 illustrated the excellent performance of RetiZero in identifying 15 fundus diseases through image-to-image retrieval. The overall scores for Top1, Top3, and Top5 are 0.8537, 0.9279, and 0.9500, respectively, representing an improvement of 9.35%, 4.79%, and 3.22% over RETFound [14], and 300.22%, 121.11%, 74.00% over FLAIR [16]. In addition, RetiZero demonstrated the best performance across all categories compared to RETFound [14] and FLAIR [16] (Supplementary Figure 3). Moreover, in the more challenging Eye-52 dataset, RetiZero achieved overall Top1, Top3, and Top5 scores of 0.7257, 0.8432, and 0.8860, respectively (Figure 2 (e)). Improved by 12.35%, 7.85%, and 6.30% over RETFound [14], and 767.75%, 389.88%, 271.51% over FLAIR [16], respectively. Furthermore, in analysis of individual disease, RetiZero demonstrated great potential, particularly in identifying several rare fundus diseases such as Bietti crystalline dystrophy (Top1: 0.8605, Top3: 0.9360, and Top5: 0.9419), chorioretinal coloboma (Top1: 0.8192, Top3: 0.8927, and Top5: 0.9096), and punctate inner choroidopathy multifocal choroiditis (Top1: 0.9022, Top3: 0.9464, and Top5: 0.9621) (Supplementary Figure 4). More details on the 52 disease categories can be found in Supplementary Figure 4. In addition, we also calculated Precision@1, Precision@3, and Precision@5 to comprehensively evaluate RetiZero’s performance in the task of fundus disease identification through image-to-image retrieval. Figures 2 (d) and Figures 2 (e) demonstrate that RetiZero achieved the highest Precision@1, Precision@3, and Precision@5 on both testing datasets, EYE-15 and EYE-52. Meanwhile, RetiZero demonstrated the best performance across most categories compared to RETFound [14] and FLAIR [16] (Supplementary Figure 5 and Supplementary Figure 6). Moreover, Figure 2 (f) shows an example of the Top5 prediction results from RetiZero, RETFound, and FLAIR. It can be seen that RetiZero achieved superior retrieval performance compared to RETFound and FLAIR, further demonstrating its excellent feature representation capabilities. Furthermore, Supplementary Figure 7 presents heatmaps illustrating the weights of different foundational models for various fundus diseases. RetiZero’s weights were more precisely concentrated on the regions affected by different fundus diseases. This precise focus underscores RetiZero’s capability to accurately identify a range of fundus diseases, including rare ones, providing significant evidence of its diagnostic proficiency.
2.3 Clinical evaluation by ophthalmology experts from different countries
To comprehensively evaluate RetiZero’s capability in fundus disease recognition without retraining the model, we randomly selected two samples from each category of the EYE-52 dataset, creating a new subset called EYE52-sub with a total of 104 instances. 14 ophthalmic experts from Singapore, the United States, and China were invited to diagnose the 104 samples in the EYE52-sub dataset. Among them, four ophthalmic experts have 3 to 5 years of clinical experience, six have 5 to 10 years of clinical experience, and four have more than 10 years. Specifically, we developed an online fundus image reading system and uploaded the 104 samples to the server, as shown in Figure 3. Mimicking the zero-shot setup, we provided 52 disease options on the webpage as prompts. During the image reading process, the clinicians selected diagnostic results from the 52 disease categories based on the image content. Finally, each expert was asked to assess their confidence in their diagnostic results. As shown in Figure 3, the diagnostic accuracy of the 14 ophthalmologists ranges from 0.365 to 0.788, and the median values between ophthalmologists is 0.582, while RetiZero’s zero-shot Top1, Top3, and Top5 accuracies are 0.308, 0.635, and 0.798, respectively. Therefore, RetiZero’s zero-shot Top3 performance is comparable to that of most ophthalmologists, and its Top5 performance surpasses that of all ophthalmologists. Furthermore,the fundus disease identification performance by image-to-image retrieval for RetiZero achieved Top1, Top3, and Top5 accuracies of 0.6837, 0.7449, and 0.7959, respectively. Therefore, RetiZero’s Top1 accuracy in identifying fundus diseases through image-to-image retrieval surpasses that of most ophthalmologists. These experimental results further demonstrate that RetiZero can achieve performance comparable to experienced ophthalmologists through zero-shot and image retrieval methods without retraining the model.
Figure 2: Overall Top1, Top3, and Top5 scores for zero-shot based fundus disease recognition and Fundus diseases identification by image-to-image retrieval. (a) The zero-shot performance on the EYE-15 dataset, which contains 30,089 fundus images including 14 common fundus diseases and a normal condition. (b) The zero-shot performance on the EYE-52 dataset, which contains 7,007 fundus images including 51 categories of fundus diseases and a normal condition. (c) Zero-shot fundus disease identification samples. (d) Image-to-image retrieval performance on EYE-15 dataset. (e) Image-to-image retrieval performance on the EYE-52 dataset. (f) Image-to-image retrieval samples. Supplementary Table 3 provides the incidence/prevalence of each category in the EYE-52 dataset.
2.4 Internal domain retinal disease identification
We collected three independent datasets from 5 ophthalmic clinics, named H1, H2, and H3, to validate the performance of RetiZero in internal domain retinal disease identification tasks. Supplementary Figure 11 provides the data collection process and annotation details of three datasets. "Internal domain" means that we fine-tuned and validated the model separately within each of the three datasets. The details about three datasets as shown in Supplementary Table 4 to Supplementary Table 6. As shown in Figure 4 (a), RetiZero achieved average AUCs of 0.9972, 0.9796, and 0.9930 on the three datasets, respectively, each encompassing 15, 13, and 12 different categories of retinal diseases/normal condition, respectively. These results represent improvements of 1.90%, 6.94%, and 2.85% compared to RETFound [14], 1.06%, 7.97%, and 3.02% compared to FLAIR [16]. This is particularly evident for certain retinal diseases with ambiguous features, such as macular hole, epiretinal membrane, and retinal artery occlusion. RetiZero exhibited statistically significant improvement in the performance of identifying retinal diseases, demonstrating an obvious superiority over RETFound [14] and FLAIR [16] (all p<0.05 in all of the three datasets, Figure 4 (a)).
Figure 3: Clinical evaluation. (a) Ophthalmologist diagnostic results, top1, top3 and top5 performance for zero-shot and image-to-image retrieval. (b) Our developed online fundus image reading system.
Figure 4: The receiver operating characteristic (ROC) curves for internal domain retinal diseases identification and few-shot learning.
Figure 5: Cross-domain performance of different foundation models for fundus diseases screening. Column (a) Internal evaluation: Different foundation models were adapted to each dataset by fine-tuning and internally evaluated on hold-out testing data. Columns (b) and (c) Performance on external validation sets: The three foundation models were tested on the other two external validation datasets. The disease categories and dataset strategy information are listed in Supplementary Tables 10 to 12.
2.5 Few-shot fine-tuning
Limited annotated data has consistently hindered the advancement of AI algorithms for medical image recognition. To address the challenge, we fine-tuned the model using only five samples from each fundus disease to evaluate RetiZero’s performance in identifying fundus diseases with very limited training data. The data details were provided in Supplementary Table 7 to Supplementary Table 9. As shown in Figure 4 (b), RetiZero achieved the highest AUROC scores across the three datasets compared to RETFound [14] and FLAIR [16]. In the task of identifying 15, 13, and 12 types of fundus diseases in the H1, H2, H3 dataset, RetiZero achieved AUROC values of 0.9668, 0.8585, and 0.9422 respectively, representing improvements of 7.21% to 35.10% over RETFound [14] and FLAIR [16] (all $P$-value < 0.01). These experimental results indicate that RetiZero possesses superior fundus feature representation capabilities. Even with limited annotated data samples, it can effectively learn the characteristic information of different fundus diseases in fundus images.
2.6 Cross domain fundus disease identification
To validate the robustness of RetiZero in the task of cross-domain fundus disease identification, we further reorganized the three datasets of H1, H2, and H3 and only used the data with shared categories across the three datasets. Then, we sequentially used the reorganized datasets of rH1, rH2, and rH3 as internal datasets and utilized the remaining two datasets as external testing sets to verify the robustness of different foundation models. The data information for different experimental strategies is presented in Supplementary Table 10 to 12. As shown in Figure 5, RetiZero achieved promising performance in all validation settings. Specifically, in the internal test set of the three datasets, RetiZero achieved AUROC values of 0.9984, 0.9857, and 0.9901, respectively, representing improvements of 3.12%, 8.60%, and 6.32% over RETFound (all p-values < 0.01) [14]; 0.73%, 4.38%, and 4.97% over FLAIR (all $P$-values < 0.05, Figure 5 (a)[16], with significant performance improvements observed in all testing set. In external tests, the performance of RetiZero remained similar to the internal test, with all AUROC >= 0.9124 and significantly outperformed RETFound [14] and FLAIR [16] in all tasks (all $P$-value <= 0.02, Figure 5 (b) and Figure 5 (c)). Additionally, as shown in Supplementary Figure 8 to Supplementary Figure 10, RetiZero exhibits different AUC scores in identifying various categories of retinal diseases on distinct datasets. Notably, RetiZero achieved outstanding performance in the identification of retinal diseases across most of the categories, especially in diseases with ambiguous pathologic features such as epiretinal membrane (ERM), retinal artery occlusion (RAO), and central serous chorioretinopathy (CSCR).
3 Discussion
In this study, we trained a vision-language-foundation model RetiZero for retinal imaging using vast fundus images paired with text description. It has demonstrated its strong capability in representing retinal disease features across a wide range of downstream tasks of retinal disease identification, including internal domain and cross-domain classification, few-shot fine-tuning, zero-shot recognition, and image-to-image retrieval. The performance of RetiZero is superior to two state-of-the-art LFMs, RETFound [15] and FLAIR [16]. These results collectively demonstrated the superior generalizable and robust performance of RetiZero in both common and rare retinal disease identification.
The superiority of RetiZero over RETFound [14] and FLAIR [16] can be attributed to its unique design and diverse data used for pre-training. Although the RETFound model [14], pre-trained on a large number of fundus images using the MAE architecture, can enhance the performance of various downstream tasks, it included a limited number of fundus disease categories, particularly rare fundus diseases. In addition, it lacks the incorporation of textual information, resulting in insufficient characterization of image feature attributes, making it unsuitable for text prompt-based zero-shot fundus disease screening tasks and limiting its application in clinical practice scenarios, especially for the identification of rare fundus diseases. In contrast, FLAIR [16], based on the CLIP architecture, incorporates textual description information during network training to enhance the representation of image feature attributes. However, it is pre-trained on a very limited dataset of fundus disease knowledge, leading to poor performance in zero-shot recognition tasks for rare fundus diseases. Furthermore, FLAIR lacks guidance for learning information such as lesion contours and topological structures in images, resulting in low performance in fundus disease identification through image-to-image retrieval.
To address these limitations, we developed this fundus contrastive language-image foundation model, RetiZero, which integrates the strengths of MAE self-supervised learning and CLIP contrastive learning architectures. To further enhance the model’s understanding of fundus diseases, we curated a dataset of image-text pairs covering over 400 fundus diseases, sourced from publicly available datasets, ophthalmic textbooks, and online resources for pretraining RetiZero. As a result, RetiZero is now a foundational model with extensive and comprehensive knowledge in ophthalmology.
Classification of retinal photographs for fundus disease identification is a well-studied task. Driven by comprehensive ophthalmic knowledge and contextual information from fundus images, RetiZero provides strong feature representation capability and robustness, enabling its superior performance for fundus disease identification across internal domain, cross-domain, and few-shot learning. Although the foundation model can reduce the sample size needed for training, classification tasks still require a certain number of images for fine-tuning and testing. It would be very challenging to collect enough sample sizes for rare diseases. To address this issue, we introduced image-to-image retrieval and zero-shot recognition tasks. Both tasks do not require data for fine-tuning, making them particularly useful for rare diseases. Image-to-image retrieval involves determining the category of a query fundus image based on feature similarity scores between the query and candidate images. RetiZero leverages its excellent image content representation capabilities of the MAE architecture and the textual feature alignment characteristics of the CLIP architecture. Therefore, it achieved superior performance in fundus disease retrieval tasks based on retinal image content. While the zero-shot learning setting allows for the recognition of rare diseases with as few as one sample. RetiZero learned textual knowledge of over 400 types of fundus diseases, enabling it to perform promisingly in zero-shot fundus disease recognition tasks, a feat not achievable by RETFound.
We also recognized limitations and the need for improvements in the current work. Although our collected dataset includes knowledge about over 400 types of fundus diseases, the imbalance across different categories may limit RetiZero’s performance in downstream tasks. Therefore, further enriching the dataset with varied quantities of different categories of fundus diseases, especially for rare fundus diseases, will be part of our future work. In addition, while RetiZero has shown promising performance across multiple tasks and datasets, specialized models optimized for specific tasks may outperform generic models. Therefore, we will further explore improvements of RetiZero for specific tasks.
In conclusion, the proposed feature-calibrated retinal vision-language foundation model (RetiZero) with knowledge of over 400 retinal diseases can effectively represent the rich contextual feature information in fundus images, as well effectively learn the alignment between retinal image features and textual descriptions. RetiZero achieved superior performance on feature representation and generalizability across different retinal disease recognition tasks at various ophthalmic centers, different degrees of domain drift, and very limited training samples. Particularly, the outstanding performance of RetiZero in zero-shot fundus disease identification and image-to-image retrieval-based fundus disease recognition holds significant importance for screening fundus diseases in clinical practice, especially for rare fundus diseases.
4 Methods
4.1 Dataset
4.1.1 Data for pretraining:
We utilized RETFound [14], pre-trained on over 900,000 fundus images using the MAE architecture, as the pre-trained backbone for the Image Encoder of RetiZero. Meanwhile, we introduced low-rank learnable factors into the pre-trained RETFound and leveraged the CLIP architecture to learn image-text knowledge, aiming to enhance the model’s understanding of image-text correlations and improve its feature representation capabilities. We pre-trained RetiZero using our collected dataset comprising 341,896 image-text pairs, covering over 400 fundus diseases. Since the dataset used for pre-training with the MAE architecture has been previously described in RETFound [14], this paper focuses on detailing the 341,896 image-text dataset we collected. As shown in Supplementary Table 13, the image-text pretraining data mainly consists of three parts: publicly available dataset with category information, data from the ophthalmic books with description information, and data from online resources with description. Specifically, we collected a total of 303,129 fundus images from 29 publicly available datasets, covering over 100 different categories of retinal diseases. We used these category labels as textual descriptions corresponding to the fundus images input into RetiZero. To enable RetiZero to acquire a more comprehensive knowledge of ophthalmology, we invited 10 ophthalmologists to further collect 23,228 fundus images with corresponding textual descriptions from 180 ophthalmic books. As shown in Supplementary Table 14, these images cover 414 ophthalmic descriptive labels, encompassing nearly all known fundus diseases to date. Furthermore, we also collected 28,800 fundus data with relevant descriptions from the online resources. We assembled a team of 12 ophthalmologists to manually clean and organize 15,544 images along with their corresponding textual descriptions. In summary, the dataset for pretraining RetiZero covers almost all currently known fundus diseases, integrating very comprehensive ophthalmic knowledge. We pre-trained RetiZero on the public platform Pytorch and Nvidia Geforce DGX A100 GPU (80G). The batch size was set to 128. Adam was adopted as the optimizer to optimize RetiZero. The procedure of data collection for RetiZero pretraining was provided in Figure 1.
4.1.2 Data for internal domain retinal diseases identification:
To verify the performance of the proposed RetiZero in the task of retinal disease identification, we built three datasets across multiple ophthalmic centers: health dataset 1 (H1), health dataset 2 (H2), and health dataset 3 (H3). This study was approved by the Joint Shantou International Eye Center Institutional Review Board and adhered to the principles of the Declaration of Helsinki. The data has been de-identified. In accordance with IRB regulations, if the data does not contain any identifiable patient information, informed consent is not required. As a result, this study has been granted approval to waive the need for informed consent. The clinical assessment and labeling procedure are shown in Supplementary Figure 11. H1 dataset, consists 11,414 fundus images covering 14 categories of retinal diseases and a normal condition, was collected from multiple eye clinics using different fundus cameras. We further divided H1 into training (6,942), validation (2,284), and testing (2,288) for model fine-tuning, model selection, and performance verification, respectively. More details of data information are given in Supplementary Table 4. H2 dataset consists of 7,812 fundus images, including 12 types of retinal diseases and 1 normal condition. The category and data information are given in Supplementary Table 5. To validate the performance of fine-tuning RetiZero for retinal disease identification on the H2 dataset, we partitioned the H2 dataset into training (4,682), validation (1,561), and testing (1,569) sets, respectively, for model fine-tuning, model selection, and performance evaluation. Supplementary Table 6 provides the category and data distribution information for the H3 dataset, which comprises 10,863 fundus images across 12 categories. We divided the H3 dataset into training (6,511), validation (2,174), and testing (2,178) sets for model fine-tuning, selection, and performance evaluation. In this paper, we fine-tuned RetiZero to the task of internal domain retinal disease identification on the public platform Pytorch and Nvidia Geforce 3090 GPU (24). Adam optimizer and cross-entropy loss function were adopted to guide the model fine-tuning. The total iteration epoch and batch size were set to 100 and 64, respectively.
4.1.3 Data for few-shot fine-tuning:
To evaluate the performance of RetiZero in the few-shot fine-tuning downstream task, we further reorganized the H1, H2, and H3 datasets. Specifically, we randomly selected 5 samples from each category of H1 training set, H2 training set, and H3 training set for few-shot fine-tuning, while retaining the validation and testing datasets for model selection and performance evaluation. More details about category and data distribution information are given in Supplementary Table 7 to Supplementary Table 9. In this experiment, RetiZero was fine-tuned on the public platform Pytorch and Nvidia Geforce 3090 GPU (24G). Adam optimizer and cross-entropy loss function were adopted to guide the model optimization. The total iteration epoch and batch size were set to 1000 and 32, respectively.
4.1.4 Data for cross-domain fundus disease identification:
To verify the generality and robustness of RetiZero in the task of cross-domain fundus disease identification, we invited professional doctors to re-organize the H1, H2, and H3 datasets. Ultimately, 11 overlapping categories were identified across the three datasets, which were then renamed as rH1(10,304 fundus images), rH2 (6,829 fundus images), and rH3 (10,485 fundus images). As shown in Supplementary Tables 10 to 12, we conducted three experimental settings to validate the generality and robustness of RetiZero. Specifically, we sequentially adopted rH1, rH2, and rH3 as internal datasets for model fine-tuning, selection, and internal testing, while utilizing the remaining two datasets as external testing sets to assess the generality and robustness of RetiZero. Therefore, to deploy these experimental settings, we fine-tuned RetiZero on the public platform Pytorch and Nvidia Geforce 3090 GPUs (24G). We used the Adam optimizer and cross-entropy loss function to guide the model fine-tuning. The total iteration epochs and batch size were set to 100 and 64, respectively.
4.1.5 Data for the tasks of zero-shot fundus disease recognition and fundus disease identification by image-to-image retrieval:
We combined three datasets from different hospitals, H1, H2, and H3, into a dataset named EYE-15, containing 30,089 fundus images that include 14 common fundus diseases and 1 normal category. This dataset was used to validate RetiZero’s performance in screening common fundus diseases in zero-shot and image-to-image retrieval approaches. The data distribution of each category in EYE-15 was provided in Supplementary Table 1. Moreover, we further collaborated with several ophthalmologists from multiple eye clinics to collect 7,007 fundus images by different fundus cameras (EYE-52 dataset), comprising 51 fundus diseases and 1 normal condition, to validate the performance of zero-shot fundus disease recognition and fundus disease identification by image-to-image retrieval in a more challenging setting. The data distribution was shown in Supplementary Table 2. As shown in Supplementary Table 2, EYE-52 comprises many clinically particularly rare fundus diseases, such as albinism, Bietti crystalline dystrophy, choroidal coloboma, and choroidal neoplasm. We adopted Top1, Top3, and Top 5 accuracy to evaluate the performance of RetiZero in both tasks of zero-shot fundus disease recognition and fundus disease identification by image-to-image retrieval. And, Supplementary Figure 11 provides the process of the collection for the EYE-15 and EYE-52 datasets. In this paper, we also adopted Precsion@1, Precesion@3, and Precsion@5 as the metrics to evaluate the performance of different foundation models in the task of fundus disease retrieval. Precision@N is a metric used to evaluate the performance of information retrieval systems and ranking algorithms. It is specifically used to measure the precision of the top $N$ results returned by a system. Here is the formula and its explanation:
$Precision@N=\frac{\left|R_{all}\cap R_{Re}@N\right|}{N},$ (1)
where $R_{all}$ is the set of all relevant samples for the given query, $R_{Re}@N$ represents the set of the top $N$ samples retrieved by the system in response to the query. $\left|R_{all}\cap R_{Re}@N\right|$ denotes the number of relevant samples in the top $N$ retrieved documents, that is the count of samples that are both relevant and retrieved within the top $N$ results, while $N$ is the number of top samples considered for the calculation.
4.2 Framework of RetiZero
Figure 1 provides an overview of the RetiZero framework. RetiZero integrates the advantages of MAE self-supervised learning and CLIP contrastive learning architectures. Specifically, the model is built upon the MAE-based pre-trained backbone network RETFound [14], whose weights are frozen to preserve the model’s representation capability for complex semantic information such as lesion contours and topological structures in retinal images. Meanwhile, we introduced low-rank learnable factors into the pre-trained RETFound and leveraged the CLIP architecture to learn image-text knowledge, aiming to enhance the model’s understanding of image-text correlations and improve its feature representation capabilities. Furthermore, we incorporated an uncertainty vision-language feature calibration method based on Dirichlet reparameterization into the contrastive vision-language pretraining framework to further refine visual-language features in the high-dimensional embedding space, thereby enhancing the model’s ability to represent complex features in fundus images. Ultimately, RetiZero is obtained, which integrates the advantages of both MAE and CLIP architectures, providing feature support for subsequent downstream tasks. We will introduce the components of RetiZero in detail in the following sections.
4.2.1 Image Encoder:
As shown in Figure 1, the image encoder consists of MAE-based SSL pre-trained backbone and low-rank learnable factors. MAE is a widely used self-supervised learning approach that employs a simple autoencoder approach to reconstruct the original signal based on partial observations. MAE-based SSL pretraining can guide the network to focus on the rich structural information and contextual features in the images. Therefore, RETFound [14], pre-trained on over 900,000 fundus images, is adopted as our MAE-based pre-trained backbone. Low-rank learnable factors (LoRA) are a parameter-efficient transfer learning method based on reparameterization [17], which utilizes low-rank representations to minimize the number of trainable parameters. It enables a pre-trained large foundation model to incorporate new knowledge into new target tasks, demonstrating robust and state-of-the-art (SOTA) performance in various parameter-efficient transfer learning tasks. Therefore, we utilize low-rank learnable factors to introduce retinal feature description information into the image encoder of RetiZero, enhancing its capacity to represent feature attributes of retinal images. Specifically, given the input token sequence $F_{in}\in R^{B\times N\times C_{in}}$ and the output token sequence $F_{out}\in R^{B\times N\times C_{out}}$ obtained by the projection layer $W\in R^{C_{out}\times C_{in}}$, LoRA assumes that updates to $W$ should be gradual and stable. Therefore, we apply low-rank approximations to delineate this gradual update. First, freeze the transformer layer to keep W fixed while adding a bypass to complete the low-rank approximation. And, the bypass consists of two linear mapping layers, $A\in R^{r\times C_{in}}$ and $B\in R^{C_{out}\times r}$, where $r\ll\left\{C_{in},C_{out}\right\}$. Thus, the processing of the update layer $\hat{W}$ can be described as:
$F_{out}=\hat{W}F_{in},$ (2)
$\hat{W}=W+\nabla W=W+BA.$ (3)
Since the multi-head self-attention mechanism determines which regions to attend based on cosine similarity, LoRA was applied to the projection layers of query, key, or value to influence the attention scores. Therefore, we apply LoRA to the query and value projection layers for low-rank approximation optimization, thus the processing strategy for multi-head self-attention becomes:
$Att\left(Q,K,V\right)=Softmax\left(\frac{QK^{T}}{\sqrt{C_{out}}}+B\right)V,$ (4)
$Q=\hat{W}_{q}F=W_{q}+B_{q}A_{q}F,$ (5)
$K=W_{k}F,$ (6)
$V=\hat{W}_{v}F=W_{v}F+B_{v}A_{v}F,$ (7)
where $W_{q}$, $W_{k}$, and $W_{v}$ are frozen projection layers of RETFound, while $A_{q}$, $B_{q}$, $A_{v}$, and $B_{v}$ are trainable LORA factors.
4.2.2 Text Encoder:
Descriptions of fundus images are typically more challenging than those of natural images, as they often contain numerous specialized clinical medical terms, sometimes even comprising multiple lesion signs or sentences. Therefore, in this paper, we utilize the BioClinicalBERT [18] model pre-trained on medical texts from the MIMIC III dataset as the text encoder to obtain clinically-aware textual embeddings.
4.3 Uncertainty-based feature calibration for guiding RetiZero pretraining
In this paper, we further introduced an uncertainty vision-language feature calibration method based on Dirichlet reparameterization [19, 20] into the contrastive vision-language pretraining framework, to further calibrate visual-language features in the high-dimensional embedding space for enhancing the robustness of the model to represent complex features in fundus images. Specifically, as shown in Figure 1, RetiZero’s pretraining consists of a fundus image encoder and a text encoder. The linear layer serves as a projection head for both the image encoder and the text encoder, mapping the acquired features to a 512-dimensional embedding feature space. Let assume $\phi=\left\{\phi_{E},\phi_{H}\right\}$ denotes image encoder ($\phi_{E}$) and corresponding projection head ($\phi_{H}$). Given a fundus image $X_{i}$, the image encoder is adopted to obtain feature representation of $F_{Img}=\phi_{E}\left(X_{i}\right)$. Meanwhile, $\psi=\left\{\psi_{E},\psi_{H}\right\}$ is used to represent text encoder ($\psi_{E}$) and corresponding projection head ($\psi_{H}$). The text encoder ($\psi_{E}$) is adopted to extract feature embedding $F_{T}=\psi_{E}(X_{T})$ from text input $X_{T}$.Then, image projection head ($\phi_{H}$) and text projection head ($\psi_{H}$) are utilized to map the independent modality representations into a joint unit hyper-sphere space: $I=\frac{\phi_{H}\left(F_{Img}\right)}{\|\phi_{H}\left(F_{Img}\right)\|}$ and $T=\frac{\psi_{H}\left(F_{T}\right)}{\parallel\psi_{H}\left(F_{T}\right)\parallel}$, respectively. The similarity between the input image ($X_{i}$) and input text ($X_{T}$) are evaluated by the cosine similarity based on the normalized features: $I^{Tr}T_{T}$, where $Tr$ represents the transpose operator. With obtained similarity metrics, the optimization goal of the contrastive-based learning pre-training approach is to minimize the distance of features between paired images and text descriptions while maximizing the distance between features of unpaired samples. Specifically, assuming that a batch contains $N$ samples, $I_{i}\in\left\{I_{1},I_{2},...,I_{N}\right\}$ and $T_{i}\in\left\{T_{1},T_{2},...,T_{N}\right\}$, represent image feature vector and text feature vector of each sample, while and $G=\left\{0,1,...,N-1\right\}$ is the corresponding category label, respectively. To guide model optimization, we use the following loss function.
$L_{Con}=L_{Em}+L_{Dl},$ (8)
$L_{Em}=\frac{1}{2}\left(\sum\nolimits^{N}_{i=1}-log\left(\frac{exp\left(I^{Tr}_{i}T_{i}\right)}{\sum\nolimits^{N}_{k-1}exp\left(I^{Tr}_{i}T_{k}\right)}\right)+\sum\nolimits^{N}_{i=1}-log\left(\frac{exp\left(T^{Tr}_{i}I_{i}\right)}{\sum\nolimits^{N}_{k-1}exp\left(T^{Tr}_{i}I_{k}\right)}\right)\right),$ (9)
$L_{Dl}$ is a loss function based on feature vectors which are reparametrized from similarity measures using the Dirichlet distribution. The specific implementation is as follows:
Step (1): Obtaining the evidence feature $E_{I2T}$ and $E_{T2I}$ by applying the Softplus activation function to similarity metrics between image and text feature embedding to ensure the feature values are larger than 0:
$E_{I2T}=Softplus(I^{Tr}T),andE_{T2I}=Softplus(T^{Tr}I),$ (10)
where I2T and T2I indicate image-to-text and text-to-image contrastive direction.
Step (2): Parameterizing $E_{I2T}$ and $E_{T2I}$ to Dirichlet distribution, as:
$\bm{\alpha_{I2T,k}}=E_{I2T,k}+1,\ i.e.,\alpha_{I2T,k}=e_{I2T,k}+1,e_{I2T,k}=\left\{Softmax\left(I^{Tr}_{k}T_{1}\right),...,Softmax\left(I^{Tr}_{k}T_{N}\right)\right\},$ (11)
$\bm{\alpha_{T2I,k}}=E_{T2I,k}+1,\ i.e.,\ \alpha_{T2I,k}=e_{T2I,k}+1,e_{T2I,k}=\left\{Softmax\left(T^{Tr}_{k}I_{1}\right),...,Softmax\left(T^{Tr}_{k}I_{N}\right)\right\},$ (12)
where $\bm{\alpha_{I2T,k}}$, $\bm{\alpha_{T2I,k}}$, $e_{I2T,k}$, and $e_{T2I,k}$ are the $k$-th contrastive similarity Dirichlet distribution parameters and evidence for the image-text contrastive similarity of the $k$-th sample in a batch of $N$ samples.
Step (3): Calculating the belief masses and corresponding uncertainty score as:
$b_{I2T,k}=\frac{e_{I2T,k}}{S_{I2T}}=\frac{\alpha_{I2T,k}-1}{S_{I2T}},\ u_{I2T}=\frac{N}{S_{I2T}},$ (13)
$b_{T2I,k}=\frac{e_{T2I,k}}{S_{T2I}}=\frac{\alpha_{T2I,k}-1}{S_{T2I}},\ u_{T2I}=\frac{N}{S_{T2I}},$ (14)
where $S_{I2T}=\sum\nolimits^{N}_{k=1}\left(e_{I2T,k}+1\right)=\sum\nolimits^{N}_{k=1}\alpha_{I2T,k}$ and $S_{T2I}=\sum\nolimits^{N}_{k=1}\left(e_{T2I,k}+1\right)=\sum\nolimits^{N}_{k=1}\alpha_{T2I,k}$ are the Dirichlet intensities of image-to-text and text-to-image, respectively, used to constrain $1=\sum\nolimits^{N}_{k=1}b_{I2T,k}+u_{I2T}$ and $1=\sum\nolimits^{N}_{k=1}b_{T2I,k}+u_{T2I}$ It can be seen from Eq. 13, and Eq. 14 the probability assigned to $k$-th sample is proportional to the observed similarity evidence for sample $k$. Conversely, if less total similarity evidence was obtained, the greater the total uncertainty.
In this study, we associate the Dirichlet distribution with the distribution of feature similarity between images and text descriptions, thereby obtaining belief masses and corresponding overall uncertainty score for the similarity of images and text description for each sample of a batch, based on the evidence collected from the feature similarity matrix. Therefore, we could work out the Dirichlet distribution parameter of $\bm{\alpha_{I2T}}=\left[{}\alpha_{I2T,1},...,\alpha_{I2T,N}\right]$ and $\alpha_{T2I}=\left[{}\alpha_{T2I,1},...,\alpha_{T2I,N}\right]$ for image-to-text, and text-to-image, while obtaining the multinomial opinions $D\left(p_{I2T,i}|\alpha_{I2T,i}\right)$ and $D\left(p_{T2I,i}|\alpha_{T2I,i}\right)$, where $p_{I2T,i}$ and $p_{T2I,i}$ were the sample assignment probabilities on a simplex. Therefore, the loss function for the reparameterized similarity matrix as follows:
$L_{Dl}=L^{I2T}_{Dl}+L^{T2I}_{Dl},$ (15)
where,
$L^{I2T}_{Dl}=L^{I2T}_{Dl-CE}+\lambda\ast L_{KL},$ (16)
$L^{T2I}_{Dl}=L^{T2I}_{Dl-CE}+\lambda\ast L_{KL},$ (17)
where $L_{Dl-CE}$ ($L^{I2T}_{Dl-CE}$ and $L^{T2I}_{Dl-CE}$) was used to ensure that the correct prediction for the sample with highest similarity between image and text yielded more evidence than other samples, while $L_{KL}$ was used to ensure that incorrect predictions would yield less evidence, and $\lambda$ was the balance factor that was gradually increased so as to prevent the model from paying too much attention to the $KL$ divergence in the initial stage of training, which might result in a lack of good exploration of the parameter space and cause the network to output a flat uniform distribution.
$L_{Dl-CE}=\int\left[\sum\nolimits^{N}_{k=1}-y_{k}log\left(p_{k}\right)\right]\frac{1}{\beta\left(\alpha_{i}\right)}\prod\nolimits^{N}_{k=1}p^{\alpha_{k}-1}_{k}dp_{k}=\sum\nolimits^{N}_{k=1}y_{k}\left(\psi\left(S_{k}\right)-\psi\left(\alpha_{k}\right)\right),$ (18)
where $\psi\left(\right)$ was the digamma function, while $\beta\left(\right)$ is the multinomial beta function for the concentration parameter $\bm{\alpha}$.
$L_{KL}=log\left(\frac{\Gamma\left(\sum\nolimits^{N}_{k=1}\hat{\alpha}_{k}\right)}{\Gamma\left(N\right)\prod\nolimits^{N}_{k=1}\Gamma\left(\sum\nolimits^{N}_{k=1}\hat{\alpha}\right)}\right)+\sum\nolimits^{N}_{k=1}\left(\hat{\alpha}_{k}-1\right)\left[\psi\left(\hat{\alpha}_{k}\right)-\psi\left(\sum\nolimits^{N}_{k-1}\hat{\alpha}_{k}\right)\right],$ (19)
where $\hat{\alpha}=y+\left(1-y\right)\odot\alpha$ is the adjusted parameter of the Dirichlet distribution which could avoid penalizing the evidence of the ground-truth class to 0, and $\Gamma\left(\right)$ is the gamma function.
In general, as shown in Eq. 9 and Eq. 15 to Eq. 19, the loss function we designed can guide the network to focus on the feature differences in image-text similarity while further improving its robustness through mapping the features of the image-text similarity matrix to the Dirichlet distribution space to guide model optimization.
4.4 Definition of Dirichlet distribution
The Dirichlet distribution was parameterized by its concentration K parameters $\bm{\alpha}=\left[\alpha_{1},...,\alpha_{k}\right]$ [20, 19]. Therefore, the probability density function of the Dirichlet distribution was computed as:
$D\left(P|\alpha\right)=\begin{cases}\frac{1}{\beta\left(\alpha\right)}\prod\nolimits^{K}_{k=1}p^{\alpha_{k}-1}_{k}&for\ P\in S_{K}\\ 0&Otherwise\end{cases},$ (20)
where $S_{K}$ was the $K$-dimensional unit simplex:
$S_{K}=\left\{P|\sum\nolimits^{K}_{k=1}p_{k}=1\right\},0\leq p_{k}\leq 1,$ (21)
where $\beta\left(\alpha\right)$ represented the $K$-dimensional multinomial beta function.
5 Code Availability
The code is available at https://github.com/LooKing9218/RetiZero.
6 Data Availability
The publicly available datasets used for pre-training are available at the following links and references:
ACRIMA [21], BEH [22], DeepDRiD [23], DR1-2 [24], E-ophta [25], AIROGS [26], DeepEyeNet [27], FIVES [28], G1020 [29], Glaucoma dataset [30, 31], IDRiD [32], JICHI [33], REFUGE [34], ORIGA [35], PARAGUAY [36], EyePACS AirDoc [37], JSIEC [8], RFMid [38]. Additional data sets supporting the findings of this study are not publicly available due to the confidentiality policy of the Chinese National Health Council and institutional patient privacy regulations. However, they are available from the corresponding authors upon request. For replication of the findings and/or further academic and AI-related research activities, data may be requested from the corresponding author H.C. (drchenhaoyu@gmail.com), and any requests will be responded to within 10 working days. Source data are provided in this paper.
Supplementary files
Table 1: Data details of EYE-15 dataset.
   Category Number
   Asteroid Hyalosis 60
   Age-related Macular Degeneration 3,291
   Central Serous Chorioretinopathy 1,715
   Diabetic Retinopathy 4,422
   Epiretinal Membrane 1,620
   Glaucoma 2,160
   Macular Hole 587
   Normal 5,281
   Pathologic Myoia 2,072
   Retinal Artery Occlusion 560
   Retinal Detachment 2,980
   Retinitis Pigmentosa 685
   Retinal Vein Occlusion 2,446
   Tessellated Fundus 1,139
   Vogt-Koyanagi-Harada 1,071
   Total 30,089
Table 2: Data details of EYE-52 dataset.
   Category Number
   Acute Posterior Multifocal Placoid Pigment Epitheliopathy 24
   Acute Retinal Necrosis 89
   Albinism 95
   Angioid Streaks 83
   Asteroid Hyalosis 150
   Behcet disease 17
   BEST disease 23
   Bietti Crystalline Dystrophy 172
   Branch Retinal Vein Occlusion 200
   Central Retinal Vein Occlusion 200
   Central Serous Chorioretinopathy 200
   Chorioretinal Coloboma 177
   Choroidal Metastasis 5
   Choroidal Rupture 118
   Commotio Retinae 80
   Moderate Non-proliferative Diabetic Retinopathy 200
   Severe Non-proliferative Diabetic Retinopathy 200
   Proliferative Diabetic Retinopathy 200
   Epiretinal Membrane 200
   Geographic Atrophy 300
   Glaucoma 300
   Hypertensive Retinopathy 200
   Intraocular Foreign Body 7
   Leber Congenital Amaurosis 31
   Leukemic Retinopathy 12
   Lipemia Retinalis 22
   Macular Hole 200
   Morning Glory Anomaly 131
   Multiple Evanescent White-Dot Syndromes 53
   Myelinated Nerve Fiber 515
   Neovascular Age-related Macular Degeneration 200
   Normal 200
   Optic Atrophy 140
   Optic Disc Coloboma 65
   Pathologic Myopia 300
   Pigmented paravenous retinochoroidal atrophy 8
   Polypoidal Choroidal Vasculopathy 300
   Presumed Ocular Histoplasmosis Syndrome 6
   Punctate Inner Choroidopathy_Multifocal Choroiditis 317
   Purtscher Retinopathy 13
   Retinal Arterial Macroaneurysm 96
   Retinal Detachment 300
   Retinal Racemose Hemangioma 6
   Retinitis Pigmentosa 300
   Retinoblastoma 22
   Roth Spots 59
   Serpiginous Choroidopathy 35
   Stargardt disease 95
   Syphilitic Uveitis 4
   Valsalva Retinopathy 92
   Vogt-Koyanagi-Harada disease 200
   X-linked Retinoschisis 45
   Total 7,007
Table 3: Incidence/Prevalence of each category in EYE-52 dataset.
Category Number Incidence /Prevalence Reference
Acute Posterior Multifocal Placoid Pigment Epitheliopathy 24 I: 0.15 /100,000 https://eyewiki.aao.org/Acute_Posterior_Multifocal_Placoid_Pigment_Epitheliopathy
Acute Retinal Necrosis 89 I: 0.063 /100,000 https://www.aao.org/eyenet/article/diagnosis-and-treatment-of-acute-retinal-necrosis
Albinism 95 P: 0.667$\sim$2 /100,000 https://eyewiki.aao.org/Albinism
Angioid Streaks 83 P: 6.5 /100,000 https://europepmc.org/article/med/37868801
Asteroid Hyalosis 150 P: 8 /1000
Behcet disease 17 P: 0.12 /100,000 https://www.uptodate.com/contents/clinical-manifestations-and-diagnosis-of-behcet-syndrome
BEST disease 23 P: 0.787 /100,000 https://www.ncbi.nlm.nih.gov/pmc/articles/PMC10910552
Bietti Crystalline Dystrophy 172 P: 1.493 /100,000 https://medlineplus.gov/genetics/condition/bietti-crystalline-dystrophy/
Branch Retinal Vein Occlusion 200 P: 4.42 /1000 https://emedicine.medscape.com/article/1223498-overview
Central Retinal Vein Occlusion 200 P: 1$\sim$2 /1000 https://www.ncbi.nlm.nih.gov/pmc/articles/PMC3178209/
Central Serous Chorioretinopathy 200 P: 9.9 /100,000 in males and 1.7/100,000 in females https://pubmed.ncbi.nlm.nih.gov/22788735
Chorioretinal Coloboma 177 P: 5$\sim$22 /100,000 https://eyewiki.aao.org/Coloboma
Choroidal Metastasis 5 P: 2.3% to 9.2% in patients with cancer https://www.e-retina.or.kr/journal/view.html?doi=10.21561/jor.2020.5.1.52
Choroidal Rupture 118 P: $\sim$10 /100,000 https://www.opticianonline.net/content/features/choroidal-rupture
Commotio Retinae 80 I: 2.6% in orbital trauma https://link.springer.com/referenceworkentry/10.1007/978-3-642-35951-4_979-1
Moderate Non-proliferative Diabetic Retinopathy 200 P: 10.6% in DR https://www.ncbi.nlm.nih.gov/pmc/articles/PMC9319242
Severe Non-proliferative Diabetic Retinopathy 200 P: 1.2% in DR https://www.ncbi.nlm.nih.gov/pmc/articles/PMC9319242
Proliferative Diabetic Retinopathy 200 P: 9.9% in DR https://www.ncbi.nlm.nih.gov/pmc/articles/PMC9319242
Epiretinal Membrane 200 P: 2$\sim$20in various groups https://emedicine.medscape.com/article/1223882-overview
Geographic Atrophy 300 P: 0.81% https://www.reviewofoptometry.com/article/sizing-up-geographic-atrophy
Glaucoma 300 P: 3.54% https://www.aaojournal.org/article/S0161-6420(14)00433-3/pdf
Hypertensive Retinopathy 200 P: 2$\sim$17in various groups https://eyewiki.aao.org/Hypertensive_Retinopathy
Intraocular Foreign Body 7 I: 0.16 /100,000 https://www.ncbi.nlm.nih.gov/books/NBK576415
Leber Congenital Amaurosis 31 P: 2$\sim$3 /100,000 https://eyewiki.aao.org/Leber_Congenital_Amaurosis
Leukemic Retinopathy 12 P: 9$\sim$90in leukemia https://www.ncbi.nlm.nih.gov/pmc/articles/PMC4687193
Lipemia Retinalis 22 P: 23% in chylomicronaemia https://en.wikipedia.org/wiki/Lipaemia_retinalis
Macular Hole 200 I: 0.33% https://emedicine.medscape.com/article/1224320-overview
Morning Glory Anomaly 131 P: 2.6 /100,000 https://onlinelibrary.wiley.com/doi/full/10.1111/aos.12778
Multiple Evanescent White-Dot Syndromes 53 I: 0.45 /100,000 https://www.reviewofophthalmology.com/article/an-update-on-white-dot-syndromes
Myelinated Nerve Fiber 515 P: 0.57% https://pubmed.ncbi.nlm.nih.gov/2338989
Neovascular Age-related Macular Degeneration 200 P: 3% in the oldest https://www.ncbi.nlm.nih.gov/pmc/articles/PMC10262804
Normal 200 unknown
Optic Atrophy 140 I: 2.86 /100,000 https://www.ncbi.nlm.nih.gov/books/NBK559130
Optic Disc Coloboma 65 P: 8.9 /100,000 https://pubmed.ncbi.nlm.nih.gov/30549247
Pathologic Myopia 300 P: 0.2-3.8% https://eyewiki.aao.org/Pathologic_Myopia_(Myopic_Degeneration)
Pigmented paravenous retinochoroidal atrophy 8 P: <0.1 /100,000 https://www.orpha.net/en/disease/detail/251295
Polypoidal Choroidal Vasculopathy 300 P: 4$\sim$9.8in presumed AMD https://eyewiki.aao.org/Polypoidal_Choroidal_Vasculopathy#Prevalence_and_Incidence:
Presumed Ocular Histoplasmosis Syndrome 6 I: 1.35 /100,000 in U.S. https://www.ncbi.nlm.nih.gov/pmc/articles/PMC9361689
Punctate Inner Choroidopathy_Multifocal Choroiditis 317 I: 0.04 /100,000 https://www.ncbi.nlm.nih.gov/pmc/articles/PMC8675391
Purtscher Retinopathy 13 P: 0.024 /100,000 https://journals.lww.com/ajg/fulltext/2022/10002/s1815_purtscher_retinopathy__a_rare_clinical.1815.aspx
Retinal Arterial Macroaneurysm 96 P: 0.22 /1000 https://www.aao.org/eyenet/article/diagnosis-of-retinal-arterial-macroaneurysm
Retinal Detachment 300 I: 7.79 /100,000 https://www.ncbi.nlm.nih.gov/pmc/articles/PMC9922621
Retinal Racemose Hemangioma 6 unknown https://www.aao.org/education/disease-review/retinal-hemangiomas
Retinitis Pigmentosa 300 P: 0.2$\sim$0.33 /1000 https://www.orpha.net/en/disease/detail/791
Retinoblastoma 22 P: 5$\sim$6.67 /100,000 https://www.ncbi.nlm.nih.gov/books/NBK1452
Roth Spots 59 P: <5% in infective endocarditis https://www.ncbi.nlm.nih.gov/pmc/articles/PMC7417078/
Serpiginous Choroidopathy 35 P: 0.2$\sim$5in uveitis https://www.ncbi.nlm.nih.gov/pmc/articles/PMC6407399
Stargardt disease 95 P: 10$\sim$12.5 /100,000 https://www.orpha.net/en/disease/detail/827
Syphilitic Uveitis 4 I: 1.25 /1000 in syphilis cohort https://www.tandfonline.com/doi/full/10.1080/22221751.2023.2290836
Valsalva Retinopathy 92 unkown https://emedicine.medscape.com/article/1228106-overview?form=fpf
Vogt-Koyanagi-Harada disease 200 P: 1%-4% in U.S. https://eyewiki.aao.org/Vogt-Koyanagi-Harada_Disease
X-linked Retinoschisis 45 P: 4$\sim$20 /100,000 in males https://www.orpha.net/en/disease/detail/792
Total 7,007
Table 4: Data distribution of H1.
Category Training Validation Testing Total
AH 36 12 12 60
AMD 865 289 289 1,443
CSCR 257 86 86 429
DR 996 332 333 1,661
ERM 180 60 61 301
Glaucoma 615 205 206 1,026
MH 159 53 53 265
Normal 1275 425 425 2,125
PM 517 173 173 863
RAO 109 37 37 183
RD 815 272 272 1,359
RP 184 61 62 307
RVO 390 130 130 650
TF 286 96 96 478
VKH 158 53 53 264
Total 6,842 2,284 2,288 11,414
AH: Asteroid Hyalosis, AMD: Age-related Macular Degeneration, CSCR: Central Serous Chorioretinopathy, DR: Diabetic Retinopathy, ERM: Epiretinal Membrane, MH: Macular Hole, PM: Pathologic Myopia, RAO: Retinal Artery Occlusion, RD: Retinal Detachment, RP: Retinitis Pigmentosa, RVO: Retinal Vein Occlusion, TF: Tessellated Fundus, VKH: Vogt-Koyanagi-Harada disease.
Table 5: Data distribution of H2.
Category Training Validation Testing Total
AMD 502 167 168 837
CSCR 320 107 107 534
DR 757 253 253 1,263
ERM 300 100 100 500
Glaucoma 409 136 137 682
MH 193 64 65 322
Normal 877 293 293 1,463
PM 138 46 47 231
RAO 56 19 19 94
RD 406 135 136 677
RVO 241 80 81 402
TF 396 132 133 661
VKH 87 29 30 146
Total 4,682 1,561 1,569 7,812
AMD: Age-related Macular Degeneration, CSCR: Central Serous Chorioretinopathy, DR: Diabetic Retinopathy, ERM: Epiretinal Membrane, MH: Macular Hole, PM: Pathologic Myoia, RAO: Retinal Artery Occlusion, RD: Retinal Detachment, RVO: Retinal Vein Occlusion, TF: Tessellated Fundus, VKH: Vogt-Koyanagi-Harada disease.
Table 6: Data distribution of H3.
Category Training Validation Testing Total
AMD 606 202 203 1,011
CSCR 451 150 151 752
DR 898 300 300 1,498
ERM 491 164 164 819
Glaucoma 271 90 91 452
Normal 1015 339 339 1,693
PM 586 196 196 978
RAO 169 57 57 283
RD 566 189 189 944
RP 226 76 76 378
RVO 836 279 279 1,394
VKH 396 132 133 661
Total 6,511 2,174 2,178 10,863
AMD: Age-related Macular Degeneration, CSCR: Central Serous Chorioretinopathy, DR: Diabetic Retinopathy, ERM: Epiretinal Membrane, PM: Pathologic Myoia, RAO: Retinal Artery Occlusion, RD: Retinal Detachment, RP: Retinitis Pigmentosa, RVO: Retinal Vein Occlusion, VKH: Vogt-Koyanagi-Harada disease.
Table 7: 5-Shot training strategy information in on H1 dataset.
Category Training Validation Testing
AH 5 12 12
AMD 5 289 289
CSCR 5 86 86
DR 5 332 333
ERM 5 60 61
Glaucoma 5 205 206
MH 5 53 53
Normal 5 425 425
PM 5 173 173
RAO 5 37 37
RD 5 272 272
RP 5 61 62
RVO 5 130 130
TF 5 96 96
VKH 5 53 53
Total 75 2,284 2,288
AH: Asteroid Hyalosis, AMD: Age-related Macular Degeneration, CSCR: Central Serous Chorioretinopathy, DR: Diabetic Retinopathy, ERM: Epiretinal Membrane, MH: Macular Hole, PM: Pathologic Myopia, RAO: Retinal Artery Occlusion, RD: Retinal Detachment, RP: Retinitis Pigmentosa, RVO: Retinal Vein Occlusion, TF: Tessellated Fundus, VKH: Vogt-Koyanagi-Harada disease.
Table 8: 5-Shot training strategy information in on H2 dataset.
Category Training Validation Testing
AMD 5 167 168
CSCR 5 107 107
DR 5 253 253
ERM 5 100 100
Glaucoma 5 136 137
MH 5 64 65
Normal 5 293 293
PM 5 46 47
RAO 5 19 19
RD 5 135 136
RVO 5 80 81
TF 5 132 133
VKH 5 29 30
Total 65 1,561 1,569
AMD: Age-related Macular Degeneration, CSCR: Central Serous Chorioretinopathy, DR: Diabetic Retinopathy, ERM: Epiretinal Membrane, MH: Macular Hole, PM: Pathologic Myoia, RAO: Retinal Artery Occlusion, RD: Retinal Detachment, RVO: Retinal Vein Occlusion, TF: Tessellated Fundus, VKH: Vogt-Koyanagi-Harada disease.
Table 9: 5-Shot training strategy information in on H3 dataset.
Category Training Validation Testing
AMD 5 202 203
CSCR 5 150 151
DR 5 300 300
ERM 5 164 164
Glaucoma 5 90 91
Normal 5 339 339
PM 5 196 196
RAO 5 57 57
RD 5 189 189
RP 5 76 76
RVO 5 279 279
VKH 5 132 133
Total 60 2,174 2,178
AMD: Age-related Macular Degeneration, CSCR: Central Serous Chorioretinopathy, DR: Diabetic Retinopathy, ERM: Epiretinal Membrane, PM: Pathologic Myoia, RAO: Retinal Artery Occlusion, RD: Retinal Detachment, RP: Retinitis Pigmentosa, RVO: Retinal Vein Occlusion, VKH: Vogt-Koyanagi-Harada disease.
Table 10: Data distribution with rH1 as internal dataset and rH2 and rH3 as external test sets.
Category rH1 internal dataset rH2 external test set rH3 external test set
Training Validation Testing
Normal 1,275 425 425 1,463 1,693
AMD 865 289 289 837 1,011
CSCR 257 86 86 534 752
DR 996 332 333 1,263 1,498
ERM 180 60 61 500 819
Glaucoma 615 205 206 682 452
PM 517 173 173 231 978
RAO 109 37 37 94 283
RD 815 272 272 677 944
RVO 390 130 130 402 1,394
VKH 158 53 53 146 661
Total 6,177 2,062 2,065 6,829 10,485
AMD age-related macular degeneration, CSCR central serous chorioretinopathy, DR diabetic retinopathy, ERM epiretinal membrane, PM pathological myopia, RAO retinal artery occlusion, RD retinal detachment, RVO retinal vein occlusion, VKH vogt-koyamagi-harada disease.
Table 11: Data distribution with rH2 as internal dataset and rH1 and rH3 as external test sets.
Category rH1 internal dataset rH2 external test set rH3 external test set
Training Validation Testing
Normal 877 293 293 2,125 1,693
AMD 502 167 168 1,443 1,011
CSCR 320 107 107 429 752
DR 757 253 253 1,661 1,498
ERM 300 100 100 301 819
Glaucoma 409 136 137 1,026 452
PM 138 46 47 863 978
RAO 56 19 19 183 283
RD 406 135 136 1,359 944
RVO 241 80 81 650 1,394
VKH 87 29 30 264 661
Total 4,093 1,365 1,371 10,304 10,485
AMD age-related macular degeneration, CSCR central serous chorioretinopathy, DR diabetic retinopathy, ERM epiretinal membrane, PM pathological myopia, RAO retinal artery occlusion, RD retinal detachment, RVO retinal vein occlusion, VKH vogt-koyamagi-harada disease.
Table 12: Data distribution with rH3 as internal dataset and rH1 and rH2 as external test sets.
Category rH1 internal dataset rH2 external test set rH3 external test set
Training Validation Testing
Normal 1,015 339 339 2,125 1,463
AMD 606 202 203 1,443 837
CSCR 451 150 151 429 534
DR 898 300 300 1,661 1,263
ERM 491 164 164 301 500
Glaucoma 271 90 91 1,026 682
PM 586 196 196 863 231
RAO 169 57 57 183 94
RD 566 189 189 1,359 677
RVO 836 279 279 650 402
VKH 396 132 133 264 146
Total 6,285 2,098 2,102 10,304 6,829
AMD age-related macular degeneration, CSCR central serous chorioretinopathy, DR diabetic retinopathy, ERM epiretinal membrane, PM pathological myopia, RAO retinal artery occlusion, RD retinal detachment, RVO retinal vein occlusion, VKH vogt-koyamagi-harada disease.
Table 13: Data details for RetiZero Pretraining.
No. Dataset Images Diseases number Source
1 ACRIMA 705 2 Diaz-Pinto A, Morales,S, Naranjo V, et al. CNNs for automatic glaucoma assessment using fundus images: an extensive validation. Biomedical engineering online, 2019, 18:1-19.
2 APTOS 3,662 5 https://www.kaggle.com/c/aptos2019-blindness-detection
3 BEH 634 2 Islam M T, Mashfu S T, Faisal A, et al.Deep learning-based glaucoma detection with cropped optic cup and disc and blood vessel segmentation. Ieee Access, 2021, 10: 2828-2841.
4 Cataract 601 >4 https://www.kaggle.com/datasets/jr2ngb/cataractdataset
5 DDR 13,673 6 https://github.com/nkicsl/DDR-dataset
6 DeepDRiD 2,256 5 Liu R, Wang X, Wu Q, et al. Deepdrid:Diabetic retinopathy—grading and image quality estimation challenge. Patterns, 2022, 3(6).
7 Diabetic Retinopathy Level Detection 4,396 5 https://www.kaggle.com/datasets/arbethi/diabetic-retinopathy-level-detection
8 Diabetic Retinopathy Organized 35,128 5 https://www.kaggle.com/datasets/dola1507108/diabetic-retinopathy-organized
9 DR1-2 1,597 7 Pires R, Jelinek H F,Wainer J, et al. Advancing bag-of-visual-words representations for lesion classification in retinal images. PloS one, 2014, 9(6): e96814.
10 DR15 34,043 5 https://www.kaggle.com/datasets/nawa393/dr15_test
11 E-ophta 463 2 Decenciere E, Cazuguel G, Zhang X, et al.TeleOphta: Machine learning and image processing methods for teleophthalmology. Irbm, 2013, 34(2): 196-203.
12 AIROGS 101,433 2 De Vente C, Vermeer K A, Jaccard N, et al. AIROGS: artificial intelligence for robust glaucoma screening challenge.IEEE transactions on medical imaging, 2023.
13 DeepEyeNet 6,048 >=13 Huang J H, Yang C H H, Liu F, et al. Deepopht: medical report generation for retinal images via deep models and visual explanation. Proceedings of the IEEE/CVF winter conference on applications of computer vision. 2021: 2442-2452.
14 FIVES 800 6 Jin K, Huang X, Zhou J, et al. Fives: A fundus image dataset for artificial Intelligence based vessel segmentation. Scientific Data, 2022, 9(1): 475.
15 G1020 1020 2 Bajwa M N, Singh G A P, Neumeier W, et al. G1020: A benchmark retinal fundus image dataset for computer-aided glaucoma detection. 2020 International Joint Conference on Neural Networks&nbsp (IJCNN). IEEE, 2020: 1-7.
16 Glaucoma dataset 364 2 Anushikha Singh, Malay Kishore Dutta, M.ParthaSarathi, Vaclav Uher and Radim Burget, “Image processing based automatic diagnosis of glaucoma using wavelet features of segmented optic disc from fundus image” Computer&amp;amp;methods and programs in biomedicine, Vol. 124, pp. 108–120, Feb. 2016, doi:10.1016/j.cmpb.2015.10.010. Ashish Issac, M.Partha Sarathi and Malay Kishore Dutta, “An adaptive threshold based image processing technique for improved glaucoma detection and classification” Computer methods and programs in biomedicine, Vol. 122 No. 2, pp. 229-244, Nov. 2015, doi: 10.1016/j.cmpb.2015.08.002.
17 IDRiD 516 5 Porwal P, Pachade S, Kokare M, et al. Idrid: Diabetic retinopathy–segmentation and grading challenge. Medical image analysis, 2020, 59: 101561.
18 JICHI 9,939 5 Takahashi H, Tampo H, Arai Y, et al. Applying artificial intelligence to disease staging: Deep learning for improved staging of diabetic retinopathy. PloS one, 2017, 12(6): e0179790.
19 Messidor 1,200 4 https://paperswithcode.com/dataset/messidor-1
20 MURED 2,208 9 https://www.kaggle.com/datasets/abhirampolisetti/multi-label-retinal-disease-mured-dataset
21 REFUGE 800 2 Orlando J I, Fu H, Breda J B, et al. Refuge challenge: A unified framework for evaluating automated methods for glaucoma assessment from fundus photographs. Medical image analysis, 2020, 59: 101570.
22 ORIGA 650 2 Zhang Z, Yin F S, Liu J, et al. Origa-light: An online retinal fundus image database for glaucoma analysis and research. 2010 Annual international conference of the IEEE engineering in medicine and biology. IEEE, 2010: 3065-3068.
23 PARAGUAY 757 7 Benítez V E C, Matto I C, Román J C M, et al. Dataset from fundus images for the study of diabetic retinopathy. Data in brief, 2021, 36: 107068.
24 Retina Dataset 601 >4 https://www.kaggle.com/datasets/jr2ngb/cataractdataset
25 Kaggle 35,126 5 https://www.kaggle.com/c/diabetic-retinopathy-detection/data
26 EyePACS AirDoc 33,978 53 Ju L, Wang X, Wang L, et al. Improving medical images classification with label noise using dual-uncertainty estimation. IEEE transactions on medical imaging, 2022, 41(6): 1533-1546.
27 JSIEC 1,000 39 Cen L P, Ji J, Lin J W, et al. Automatic detection of 39 fundus diseases and conditions in retinal photographs using deep neural networks. Nature communications, 2021, 12(1): 4828.
28 RFMid 2,531 46 Pachade S, Porwal P, Thulkar D, et al. Retinal fundus multi-disease image dataset (rfmid): A dataset for multi-disease detection research. Data, 2021, 6(2): 14.
29 ODIR5K 7,000 >=7 https://www.kaggle.com/datasets/andrewmvd/ocular-disease-recognition-odir5k
30 Ophthalmic Book 23,228 >400 180 Ophthalmic books
31 Online Resources 15,544 Online Resources
Total 341,896
Table 14: Diseases category information of the data from 180 ophthalmic books.
No. Ophthalmic labels Number
1 Achromatopsia 1
2 Acute exudative polymorphous vitelliform maculopathy 32
3 Acute idiopathic blind spot enlargement 2
4 Acute Idiopathic Maculopathy (AIM) 23
5 Acute Macular Neuroretinopathy (AMN) 25
6 Acute Posterior Multifocal Placoid Pigment Epitheliopathy (APMPPE) 128
7 Acute Retinal Necrosis (ARN) 280
8 Acute Zonal Occult Outer Retinopathy (AZOOR) 37
9 Adult-Onset Foveomacular Vitelliform Dystrophy (AFVD) 30
10 Albinism 100
11 Alport retinopathy 4
12 Amblyopia 5
13 Amyloidosis 23
14 Anemic chorioretinopathy 34
15 Angioid streaks 153
16 Annular choroidal dystrophy 13
17 Anterior Ischemic Optic Neuropathy (AION) 51
18 Arc Welder’s Maculopathy 2
19 Arteriosclerotic changes 1
20 Arteritic Anterior Ischemic Optic Neuropathy (AAION) 49
21 Asteroid hyalosis 23
22 Asteroid macular dystrophy 2
23 Autosomal Dominant Vitreoretinochoroidopathy 14
24 Behcet disease 137
25 Bergmeister papilla 18
26 Best disease 301
27 Bietti Crystalline Dystrophy (BCD) 47
28 Bilateral Diffuse Uveal Melanocytic Proliferation (BDUMP) 25
29 Birdshot chorioretinopathy 90
30 Blood–brain barrier disruption maculopathy 5
31 Blue cone monochromatism 1
32 Bothnia retinal dystrophy 8
33 Branch Retinal Atery Occlusion (BRAO) 190
34 Branch Retinal Vein Occlusion (BRVO) 326
35 Bull eye maculopathy 41
36 Cancer-associated retinopathy 2
37 Carotid-cavernous fistula 1
38 Central Areolar Choroidal Dystrophy (CACD) 57
39 Central Retinal Artery Occlusion (CRAO) 192
40 Central Retinal Vein Occlusion (CRVO) 337
41 Central Serous Chorioretinopathy (CSCR) 230
42 Cherry-red spot 34
43 Chikungunya retinitis 8
44 Chorioretinal atrophy 37
45 Chorioretinitis 60
46 Chorioretinitis sclopetaria 18
47 Choroidal coloboma 126
48 Choroidal degeneration 9
49 Choroidal detachment 60
50 Choroidal folds 81
51 Choroidal granuloma 33
52 Choroidal hemorrhage 21
53 Choroidal infarction 8
54 Choroidal infiltration 1
55 Choroidal lesion 24
56 Choroidal Neovasculization (CNV) 203
57 Choroidal nevus 165
58 Choroidal rupture 95
59 Choroidal scar 4
60 Choroidal sclerosis 3
61 Choroidal tubercle 36
62 Choroideremia 114
63 Choroiditis 96
64 Chronic granulomatous disease 3
65 Cilioretinal artery occlusion 33
66 Coats disease 208
67 Cobblestone degeneration 14
68 Combined retinal artery and vein occlusion (RAVO) 17
69 Commotio Retinae 60
70 Compressive optic neuropathy 10
71 Cone-rod dystrophy 139
72 Congenital achromatopsia 2
73 Congenital grouped retinal pigment epithelium albinotic nevi 7
74 Congenital Hypertrophy of the Retinal Pigment Epithelium (CHRPE) 208
75 Congenital optic disc pigmentation 2
76 Congenital retinal macrovessels 8
77 Contusion injury 18
78 Cotton Wool Spots (CWS) 41
79 Crystalline retinopathy 33
80 Cuticular drusen 5
81 Cystoid degeneration 3
82 Dalen-Fuchs nodule 2
83 Dark without pressure 1
84 Decompression retinopathy 7
85 Diffuse Unilateral Subacute Neuroretinitis (DUSN) 91
86 Disciform lesions 3
87 Dislocated lens 25
88 Doyne Honeycomb Retinal Dystrophy (DHRD) 84
89 Dragged vessels 5
90 Drusen 223
91 Drusenoid PED 2
92 Dry Age-related Macular Degeneration (dry AMD) 160
93 Eales disease 84
94 Endophthalmitis 52
95 Enhanced S Cone Syndrome (ESCS) 17
96 Enlarged optic cup 2
97 Epiretinal Membrane (ERM) 243
98 Erosive vitreoretinopathy 5
99 Exudative retinal detachment 35
100 Familial dominant drusen 3
101 Familial Exudative Vitreoretinopathy (FEVR) 148
102 Familial Flecked Retina syndrome 33
103 Familial internal limiting membrane dystrophy 8
104 Familial optic neuropathy 11
105 Foveal hypoplasia 10
106 Frosted-branch angiitis 51
107 Fundus Albipunctatus 42
108 Fundus Flavimaculatus 30
109 fundus heterochromia 1
110 Gas bubble 22
111 Geographic Helicoid Peripapillary Choroidopathy (GHPC) 5
112 Giant Retinal Tear (GRT) 19
113 Glaucoma 154
114 Glob perforation 1
115 Goldmann-Favre disease 9
116 Granuloma 19
117 Gyrate Atrophy 72
118 Hamartoma 19
119 Hard exudate 22
120 Heavy liquid droplet 5
121 Hemangioblastoma 34
122 High-altitude retinopathy 3
123 Hyperlipidemia 7
124 Hyperoxaluria 7
125 Hypertensive retinopathy 229
126 Hyperviscosity syndrome 3
127 Hypotony retinopathy 25
128 Idiopathic obliterative arteritis 4
129 Idiopathic sclerochoroidal calcification 7
130 Idiopathic vasculitis Aneurysms and Neuroretinitis syndrome (IRVAN) 24
131 Incontinentia pigmenti 60
132 Infection 226
133 Infection Angiostrongyliasis 2
134 Infection Calliphoridae 4
135 Infection Candidiasis 76
136 Infection Cat-Scratch Disease 78
137 Infection CMV retinitis 236
138 Infection Coccidiomycosis 9
139 Infection Cryptococcosis 6
140 Infection Cysticercosis 74
141 Infection Dengue retinopathy 21
142 Infection Filariasis 4
143 Infection Gnathostomiasis 5
144 Infection HIV retinopathy 57
145 Infection Lyme disease 3
146 Infection Malaria retinopathy 3
147 Infection Nematode 4
148 Infection Onchocerciasis 2
149 Infection Ophthalmomyiasis interna 16
150 Infection Presumed Ocular Histoplasmosis syndrome (POHS) 184
151 Infection Rift Valley fever retinitis 5
152 Infection Rubella retinopathy 17
153 Infection Syphilis 15
154 Infection syphilitic chorioretinopathy 112
155 Infection Toxocariasis 116
156 Infection Toxoplasmosis 342
157 Infection Trematode 8
158 Infection Tuberculosis 24
159 Intermediate uveitis 106
160 Intraocular Foreign Body (IOFB) 68
161 Intraocular Garamycin 9
162 Laser spots 45
163 Late-onset retinal macular degeneration (LORMD) 5
164 Lattice degeneration 113
165 Leber Congenital Amaurosis (LCA) 65
166 Leber Hereditary Optic Neuropathy (LHON) 46
167 Leber idiopathic stellate neuroretinitis 25
168 Leber miliary aneurysm 6
169 Leukemic retinopathy 157
170 Lipemia retinitis 14
171 Lipid deposition 1
172 Luetic chorioretinitis 11
173 Lupus retinopathy 76
174 Lymphoma 206
175 Macropapilla 1
176 Macular atrophy 8
177 Macular dysplasia 1
178 Macular dystrophy 44
179 Macular edema 82
180 Macular Hole (MH) 247
181 Macular infarction syndrome 1
182 Macular scar 1
183 Macular telangiectasia 60
184 Maternally inherited diabetes and deafness (MIDD) 6
185 Megalopapilla 10
186 Microcystoid degeneration 2
187 micropapilla 1
188 Microphthalmos 1
189 Mitochondrial retinal Dystrophy 7
190 Morning Glory syndrome (MGS) 83
191 Multifocal Choroiditis (MFC) 115
192 Multiple Evanescent White Dot Syndrome (MEWDS) 67
193 Myelinated Nerve Fiber (MNF) 90
194 Myopia 46
195 Nanophthalmos 1
196 Neovascular Age-related Macular Degeneration (nAMD) 320
197 Neovascularization 59
198 Neuroretinitis 72
199 Newfoundland rod–cone degeneration (NFRCD) 3
200 Nicotinic acid maculopathy 2
201 Nonarteritic Anterior Ischemic Optic Neuropathy (NAION) 72
202 Non-Proliferative Diabetic Retinopathy (NPDR) 363
203 Normal 186
204 Norrie Disease 4
205 North Carolina Macular Dystrophy (NCMD) 108
206 Ocular Ischemic Syndrome (OIS) 68
207 Ocular melanocytosis 5
208 Ophthalmic artery occlusion 8
209 Optic disc anomaly 30
210 Optic disc aplasia 3
211 Optic disc atrophy 172
212 Optic disc avulsion 29
213 Optic disc coloboma 72
214 Optic disc drusen 82
215 Optic disc dysplasia 3
216 Optic disc granuloma 11
217 Optic disc hemorrhage 11
218 Optic disc hyaline body 15
219 Optic disc hypoplasia 65
220 Optic disc metastasis 21
221 Optic disc pallor 29
222 Optic disc pit 130
223 Optic neuritis 89
224 Optic neuropathy 12
225 Outer retinal corrugation 4
226 Overlapping WDS 10
227 Panuveitis 3
228 Papilloedema 332
229 Papillomegaly 8
230 Papillophlebitis 6
231 Papillorenal syndrome 13
232 Paraneoplastic vitelliform dystrophy 2
233 Paraneoplastic-Related Retinopathy 11
234 Pathological Myopia (PM) 692
235 Pattern Dystrophy 152
236 Pearl degeneration 1
237 Peripapillary atrophy 2
238 Peripheral exudative hemorrhagic chorioretinopathy 9
239 Perivasculitis 51
240 Persistent Hyperplastic Primary Vitreous (PHPV) 60
241 Persistent Placoid Maculopathy (PPM) 13
242 Pigment Epithelial Detachment (PED) 97
243 Pigmentary retinopathy 117
244 Pigmented Paravenous Chorioretinal Atrophy (PPCRA) 9
245 Polypoidal Choroidal Vasculopathy (PCV) 76
246 Posterior microphthalmos 2
247 Posterior placoid chorioretinitis 4
248 Posterior scleritis 60
249 Posterior staphyloma 18
250 Posterior uveitis 16
251 Posterior Vitreous Detachment (PVD) 34
252 Pregnancy-associated retinopathy 53
253 Pre-retinal hemorrhage 47
254 Progressive dominantly inherited dystrophy 21
255 Progressive Outer Retinal Necrosis (PORN) 39
256 Progressive subretinal fibrosis and uveitis syndrome 8
257 Progressive systemic sclerodermic retinopathy 1
258 Proliferative Diabetic Retinopathy (PDR) 546
259 Proliferative Vitreoretinopathy (PVR) 201
260 Pseudopapilloedema 9
261 Pseudovitelliform detachment 23
262 Punctate Inner Choroidopathy (PIC) 54
263 Purtscher retinopathy 68
264 Radiation retinopathy 65
265 Reactive lymphoid hyperplasia 6
266 Relentless Placoid Chorioretinitis (RPC) 9
267 Reticular pseudodrusen 15
268 Retinal Angiomatous Proliferation (RAP) 42
269 Retinal arteriovenous malformation 11
270 Retinal artery macroaneurysm 151
271 Retinal break 15
272 Retinal cyst 2
273 Retinal Detachment (RD) 463
274 Retinal dystrophy 7
275 Retinal folds 19
276 Retinal hemorrhage 50
277 Retinal infarction 8
278 Retinal infiltrates 2
279 Retinal ischemia 9
280 Retinal pigment epitheliitis 13
281 Retinal sheathing 8
282 Retinal tear 208
283 Retinal telangiectasia 70
284 Retinal tuft 19
285 Retinitis 54
286 Retinitis Pigmentosa (RP) 302
287 Retinitis punctata albescens 13
288 Retinopathy of Prematurity (ROP) 570
289 Retinoschisis 122
290 Retrobulbar neuritis 1
291 Roth spots 15
292 RPE atrophy 13
293 RPE epithelioma 3
294 RPE hyperplasia 14
295 RPE nevus 7
296 RPE tear 8
297 Sarcoidosis 11
298 Scleritis 1
299 Sclerochoroidal calcification 7
300 Serous retinal detachment 7
301 Serpiginous choroidopathy 143
302 Shaken baby syndrome 63
303 Sickle cell retinopathy 172
304 Silicone oil 46
305 Sjögren Reticular Dystrophy 11
306 Sjögren–Larssen syndrome 2
307 Snailtrack degeneration 18
308 Snowflake degeneration 6
309 Solar/Laser Maculopathy 78
310 Sorsby Pseudoinflammatory Fundus Dystrophy (SPFD) 49
311 Staphyloma 4
312 Stargardt disease 267
313 Stationary night blindness 11
314 Submacular abscesses 1
315 Subretinal fibrosis 14
316 Subretinal fibrosis and uveitis syndrome 28
317 Subretinal hemorrhage 32
318 Sub-RPE hemorrhage 1
319 Suprachoroidal hemorrhage 2
320 Susac syndrome 11
321 Sympathetic ophthalmia 56
322 Synchysis scintillans 2
323 Systemic diseases 70
324 Takayasu retinopathy 4
325 Talc retinopathy 18
326 Terson syndrome 28
327 Tilted disc 44
328 Torpedo maculopathy 8
329 Toxicity 423
330 Toxicity Chalcosis 3
331 Toxicity Siderosis 3
332 Toxicity Tacrolimus microangiopathy 10
333 Tractional retinal detachment 1
334 Transplant-associated retinopathy 7
335 Trauma 81
336 Trauma electrocution retinopathy 3
337 Trauma gunshot 2
338 Trauma optic neuropathy 3
339 Trauma retinal pigment epitheliopathy 6
340 Trauma retinopathy 7
341 Tumor 15
342 Tumor adenocarcinoma of the RPE 3
343 Tumor adenoma of the ciliary body pigment epithelium 2
344 Tumor adenoma of the RPE 19
345 Tumor astrocytic hamartoma 4
346 Tumor cavernous hemangioma 44
347 Tumor choroidal hemangioma 182
348 Tumor choroidal melanoma 512
349 Tumor choroidal metastasis 213
350 Tumor choroidal osteoma 114
351 Tumor choroidal plasmacytoma 2
352 Tumor ciliochoroidal melanoma 29
353 Tumor combined hamartoma of the retina and RPE 117
354 Tumor congenital simple hamartoma of the RPE 2
355 Tumor leiomyoma 1
356 Tumor medulloepithelioma 6
357 Tumor melanocytoma 6
358 Tumor metastasis 4
359 Tumor myeloma 2
360 Tumor optic disc astrocytic hamartoma 5
361 Tumor optic disc astrocytoma 6
362 Tumor optic disc capillary hemangioma 15
363 Tumor optic disc glioblastoma 2
364 Tumor optic disc glioma 4
365 Tumor optic disc hemangioblastoma 5
366 Tumor optic disc hemangioma 2
367 Tumor optic disc melanocytoma 104
368 Tumor optic disc melanoma 3
369 Tumor optic disc meningioma 5
370 Tumor retinal angioma 1
371 Tumor retinal astrocytic hamartoma 111
372 Tumor retinal astrocytoma 59
373 Tumor retinal capillary hemangioma 220
374 Tumor retinal cavernous hemangioma 39
375 Tumor retinal hamartoma 2
376 Tumor retinal melanocytoma 1
377 Tumor retinal metastasis 22
378 Tumor retinal racemose hemangioma 25
379 Tumor Retinoblastoma (RB) 513
380 Tumor retinocytoma 7
381 Tumor RPE hamartomas 3
382 Tumor teratoma 1
383 Tumor uveal hemangiopericytoma 1
384 Tumor uveal melanoma 2
385 Tumor uveal metastasis 1
386 Tumor uveal schwannoma 3
387 Tumor Vasoproliferative Tumor (VPT) 48
388 Type 1 aneurysmal telangiectasis 1
389 Type 2 aneurysmal telangiectasis 3
390 Unifocal helioid choroiditis 1
391 Uveal benign reactive lymphoid hyperplasia 14
392 Uveal Effusion Syndrome (UES) 46
393 Uveitis 46
394 Valsalva retinopathy 59
395 Vascular anomaly 42
396 Vascular sheathing 4
397 Vascular tortuosity 40
398 Vasculitis 99
399 Venous stasis retinopathy 25
400 Vessel shunt 31
401 Vitreomacular Traction syndrome (VMT) 51
402 Vitreous base avulsion 3
403 Vitreous cyst 9
404 Vitreous Hemorrhage (VH) 71
405 Vitreous liquefaction 4
406 Vitreous opacity 15
407 Vitritis 25
408 Vogt-Koyanagi-Harada disease (VKH) 283
409 West African crystalline maculopathy 9
410 West Indies crinkled retinal pigment epitheliopathy 3
411 White with pressure 4
412 White without pressure 10
413 Xerophthalmia 5
414 X-linked juvenile retinoschisis (XLRS) 153
Total 23,228
Figure 1: Top1, Top3, and Top5 scores of zero-shot recognition for each category in the EYE-15 dataset.
Figure 2: Top1, Top3, and Top5 scores of zero-shot recognition for each category in the EYE-52 dataset.
Figure 3: Top1, Top3, and Top5 scores of fundus disease identification by image-to-image retrieval for each category in the EYE-15 dataset.
Figure 4: Top1, Top3, and Top 5 scores of fundus diseases identification by image-to-image retrieval for each category in the EYE-52 dataset.
Figure 5: Precision@1, Precision@3, and Precision@5 of fundus diseases identification by image-to-image retrieval for each category in the EYE-15 dataset.
Figure 6: Precision@1, Precision@3, and Precision@5 scores of fundus diseases identification by image-to-image retrieval for each category in the EYE-52 dataset.
Figure 7: Heatmaps of different foundation model weights for various fundus diseases.
Figure 8: AUC scores of different Foundation models on different datasets, where the reorganized rH1 dataset was adopted as an internal set, while both reorganized rH2 and rH3 datasets were used as external test sets.
Figure 9: AUC scores of different Foundation models on different datasets, where the reorganized rH2 dataset was adopted as an internal set, while both reorganized rH1 and rH3 datasets were used as external test sets.
Figure 10: AUC scores of different Foundation models on different datasets, where the reorganized rH3 dataset was adopted as an internal set, while both reorganized rH1 and rH2 datasets were used as external test sets.
Figure 11: For the downstream tasks, we divided the datasets as follows: H1 comprised devices 1 and 2 of JSIEC, H2 consisted of devices 1 and 2 from four other medical centers, and H3 included data from device 3 of JSIEC. Additionally, we collected data from devices D1 and D2 of all medical centers to form the EYE-15 and EYE-52 datasets, respectively.